OpenStudy (anonymous):

Given f(x) = 2x − 3 and g(x) = 4x − 4, evaluate the composite function. f [g(x)]

OpenStudy (anonymous):

g(x) = 4x − 4 f(g(x)) = f(4x − 4) = 2(4x − 4) - 3 = 8x - 8 -3 = 8x -11

OpenStudy (anonymous):

answer is 8x -11

OpenStudy (mertsj):

f(g(x)) = 2(4x-4) -3 =8x-8-3=8x-11